OpenStudy (anonymous):

evaluate the following compositions if f(x)=2x and g(x)=x+3 f(g(2)) Is the answer 2x(x+3)(2) 2x(2x+6) 4x^2+12x

OpenStudy (anonymous):

other form f(g(x))=2(x+3) f(g(2))=2(2+3) = 10